OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 4x < -16 or x + 5 greater than or equal to 15

OpenStudy (anonymous):

A number line with a closed circle on -4, shading to the left, and an open circle on 10, shading to the right. A number line with an open circle on -4, a closed circle on 10, and shading in between. A number line with a closed circle on -4, an open circle on 10, and shading in between. A number line with an open circle on -4, shading to the left, and a closed circle on 10, shading to the righ

OpenStudy (aum):

4x < -16 or x + 5 greater than or equal to 15 \(4x \lt -16\) OR \(x + 5 \ge 15\) \(x < -4\) OR \(x \ge 10\) \(x < -4\) : A number line with a open circle on -4 and shading to the left. \(x \ge 10\) : A number line with a closed circle on 10 and shading to the right.